OpenStudy (anonymous):

Use a system of equations to solve the following problem: Charles High School has a total of 94 boys and girls who play sports. If the number of boys is 22 more than twice the number of girls, how many boys play sports at this high school?

OpenStudy (anonymous):

70 48 24 80

OpenStudy (anonymous):

im out of practice with these so my math may be incorrect and its better to double check your own work rather then jsut accept my opinion however, i took the original number given and subtracted the "bonus" number of boys this gives you 94 - 22 take the answer 72 and divide by 3, for 24, the base number of girls then finish up using 24 x 2 + 22 = 70 should be the method to solve this and similar questions

hero (hero):

x = boys y = girls x + y = 94 x = 2y+22 x = 94 - y x = 2y +22 x = x 94 - y = 2y + 22 94 - 22 = 2y + y 72 = 3y 72/3 = y 24 = y 70 = x 70 boys 24 girls